OpenStudy (anonymous):

solve by substitution method 0.08x-0.01y=4.4 0.26x+0.14y=21.2

OpenStudy (anonymous):

To get rid of the decimals multiply each equation by 100 to get the equations 8x - y = 440 26x + 14y = 2120 solving the first equation for y give y=8x - 440 substituting into the second equation give 26x + 14(8x - 440) = 2120 26x + 112x - 6160 = 2120 138x = 8280 x 8280/138 = 60

OpenStudy (anonymous):

typo. i meant x = 8280/138 = 60. then substitute back into first equation to get y = 8*60-440 = 480-440 = 40